OpenStudy (anonymous):

CUT IN HALF - Conversion Math 1/4 Teaspoon baking soda! HELP!

OpenStudy (anonymous):

you want half of \(\frac{1}{4}\) in math you want \(\frac{1}{2}\times \frac{1}{4}\)